OpenStudy (jhhhhhiiiii):

Find the angle between the given vectors to the nearest tenth of a degree. u = <-5, -4>, v = <-4, -3>

OpenStudy (jhhhhhiiiii):

@mathmate

OpenStudy (jhhhhhiiiii):

@UnkleRhaukus

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

first you need the dot product, then you have to divide by the product of the absolute values then take the arc cosine

OpenStudy (jhhhhhiiiii):

I'm not sure how to do this, I'm new to it

OpenStudy (misty1212):

\[\large <x_1,y_1>\cdot<x_2,y_2>=x_1x_2+y_1y_2\]

OpenStudy (misty1212):

so in your case \[(-5)(-4)+(-4)(-3)\]

OpenStudy (jhhhhhiiiii):

Okay I got 32

OpenStudy (misty1212):

then you need the absolute values \[||<x,y>||=\sqrt{x^2+y^2}\]

OpenStudy (misty1212):

you need them for each one

OpenStudy (jhhhhhiiiii):

How do I set it up?

OpenStudy (jhhhhhiiiii):

Like the square root of -5^2+-4?

OpenStudy (misty1212):

\[\sqrt{5^2+4^2}\] for the first one

OpenStudy (jhhhhhiiiii):

-4^2*

OpenStudy (misty1212):

for the second \[\sqrt{4^2+3^2}\]

OpenStudy (misty1212):

you should get \(\sqrt{41}\) and \(5\) respectively

OpenStudy (jhhhhhiiiii):

Okay thank you

OpenStudy (misty1212):

then \[\cos(\theta)=\frac{32}{5\sqrt{41}}\]

OpenStudy (misty1212):

making \[\theta=\cos^{-1}(\frac{32}{5\sqrt{41}})\]

OpenStudy (jhhhhhiiiii):

Is it 1.78

OpenStudy (jhhhhhiiiii):

Choices: -9.1° 1.8° 0.9° 11.8°

OpenStudy (jhhhhhiiiii):

Is it 1.8?

OpenStudy (misty1212):

that is what i get, yes http://www.wolframalpha.com/input/?i=arccos(32%2F(5sqrt(41)))